# Learning Perceptually Relevant Temporal Envelope Morphing

### v3

@@ 5 Experiments
To assess the effectiveness and generalizability of our proposed audio envelope morphing method, we construct three evaluation benchmarks of increasing complexity. Each benchmark tests a different aspect of perceptual morphing: accuracy on simple synthetic rules, robustness to compositional morphs involving multiple perceptual dimensions, and performance on real-world like acoustic data. In all settings, the task is to predict the morphed envelope given two source envelopes and an interpolation weight $\alpha$, and compare it to the ground truth morph $E_{\text{morph}}$ using RMSE. We evaluate against three baselines: (1) audio mixing, which linearly interpolates between $E_{1}$ and $E_{2}$ in amplitude; (2) embed mixing, which interpolates between their latent representations and decodes the result; and (3) DTW, which performs morphing as proposed by [38] where Dynamic Time Warping algorithm [39] is used to find the optimal warping path between the two envelopes and linear interpolation is used to generate the morph.
@@ 5.3 Evaluating the generalization to naturalistic signals
While synthetic data offers interpretability and controlled supervision, it lacks the acoustic complexity and variability of real-world audio. However, fully natural morphing benchmarks lack ground-truth morphs. To balance these needs, we design a naturalistic audio envelope benchmark using real world impulse sounds arranged with synthetic structure [40] as shown in Figure 2. We extract ten diverse impulse events (e.g., a dog bark, a bicycle bell, etc.) from the Clotho dataset [41], manually isolate single impulses, and use them to generate a set of 10k tuples following the same perceptual morphing rules as described in Section 4.2). These audio clips are processed into envelope representations using the same pipeline as before.
@@ 6 Conclusion
We introduced a perceptually grounded framework for morphing temporal envelopes, informed by empirical studies on human auditory perception. By identifying key perceptual factors-such as the quantity, spacing, and placement of impulses-we codified the principles that underlie natural-sounding morphs. These principles enabled the creation of a large synthetic dataset of envelope morphing examples, which we used to train a twin network for generating perceptually valid interpolations. Our experiments demonstrate that the proposed framework consistently outperforms conventional methods across multiple benchmarks. Increasingly, modern sound synthesis and generative models are conditioned on temporal envelopes [42, 43, 44, 45]. By providing a perceptually intuitive way to morph these envelopes, our framework can enhance the controllability and naturalness of the resulting sound morphs in such systems. By bridging psychoacoustics and DL, this work advances audio generation systems toward perceptually grounded morphing - a critical capability with applications in assistive sound design, immersive media, and computational hearing research.
@@ 7 Acknowledgment
Work supported by the Sony Research Award Program (RAP). This work used Bridges2 at PSC through allocation SOC240007P from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.

### v4

@@ 5 Experiments
To assess the effectiveness and generalizability of our proposed audio envelope morphing method, we construct three evaluation benchmarks of increasing complexity. Each benchmark tests a different aspect of perceptual morphing: accuracy on simple synthetic rules, robustness to compositional morphs involving multiple perceptual dimensions, and performance on real-world like acoustic data. In all settings, the task is to predict the morphed envelope given two source envelopes and an interpolation weight $\alpha$, and compare it to the ground truth morph $E_{\text{morph}}$ using RMSE. We evaluate against three baselines: (1) audio mixing, which linearly interpolates between $E_{1}$ and $E_{2}$ in amplitude; (2) embed mixing, which interpolates between their latent representations and decodes the result; and (3) DTW, which performs morphing as proposed by [li2023morphing] where Dynamic Time Warping algorithm [muller2007dtw] is used to find the optimal warping path between the two envelopes and linear interpolation is used to generate the morph.
@@ 5.3 Evaluating the generalization to naturalistic signals
While synthetic data offers interpretability and controlled supervision, it lacks the acoustic complexity and variability of real-world audio. However, fully natural morphing benchmarks lack ground-truth morphs. To balance these needs, we design a naturalistic audio envelope benchmark using real world impulse sounds arranged with synthetic structure [xie2024audiotime] as shown in Figure 2. We extract diverse impulse events (e.g., a dog bark, a bicycle bell etc.) from the Clotho dataset [drossos2020clotho], manually isolate single impulses, and use them to generate a set of 10k tuples following the same perceptual morphing rules as described in Section 4.2). These audio clips are processed into envelope representations using the same pipeline as before.
@@ 6 Conclusion
We introduced a perceptually grounded framework for morphing temporal envelopes, guided by empirical studies on human auditory perception. By codifying key perceptual factors—such as the quantity, spacing, and placement of impulses—we built a large synthetic dataset of envelope morphing examples to train a twin network for perceptually valid interpolations. Experiments show our framework consistently outperforms conventional methods. As modern synthesis and generative models increasingly rely on temporal envelopes [chung2024t, garcia2025sketch2sound, gramaccioni2024stable, lee2024video], our approach offers a perceptually intuitive way to morph them, enhancing controllability and naturalness. Bridging psychoacoustics and deep learning, this work advances audio generation toward perceptually grounded morphing with applications in assistive sound design, immersive media, and computational hearing.
@@ 7 Acknowledgment
We thank Sony for their support of this work through the Sony Research Award Program (RAP).